# Not Like Us, Hunty: Measuring Perceptions and Behavioral Effects of Minoritized Anthropomorphic Cues in LLMs

### v1

@@ 5.4.3. Negative Perceptions of the Sociolect LLMs
These positive sentiments were contrasted by a notable subset of participants who perceived the sociolect LLM’s use of sociolect as forced or artificial. Among Queer participants, 8.2% of comments were tagged with “negative emotions,” 6.98% with “exaggerated usage,” and 5.33% with “disrespectful.” Some participants expressed frustration at the QSLM’s overuse of slang, with one noting, “Even people who use LGBTQ slang don’t talk like that constantly. It would be annoying to have an AI constantly use slang phrases.” Similarly, AAE participants criticized the AAELM, describing its language as unnatural and inauthentic. Comments such as “Agent [AAELM] using AAE sounds like a joke and not natural” were reflected in 3.41% of comments tagged as “unnatural,” 2.61% tagged as “unserious,” and 1.4% tagged as “disrespectful.” Participants also raised concerns about stereotyping and cultural insensitivity. AAE participants expressed discomfort, perceiving the AAELM’s use of AAE as stereotyping or mocking Black culture. One respondent noted, “I don’t like the idea of AI using such language, especially if it is being programmed by someone who isn’t Black.” Similarly, Queer slang participants criticized the QSLM, feeling that its use of Queer slang bordered on mockery, with comments such as, “The [QSLM] seems to be making a mockery of lgbt [sic].” These reactions highlight the risks of sociolect adaptation, especially when it involves sensitive cultural or linguistic contexts, potentially alienating users if the sociolect is perceived as inauthentic or disrespectful.
@@ 5.4.4. Positive Perceptions of the SAE-using LLMs
AAE participants often emphasized that SAELM met their expectations for how an AI should behave. Statements like “Agent [SAELM] sounds normal and like an actual AI” underscore this sentiment, with 3.8% of AAE comments tagged as “normal.” These responses suggest that SAE’s alignment with participants’ expectations of AI fosters a sense of credibility and trustworthiness. While both participant groups frequently perceived SAELM as more relatable and human-like, some responses also indicated that it could connect on a personal level. For example, 1.4% of AAE participant comments were tagged as “resonates.” Remarks such as “I have a personal connection with him” reveal that SAE, while perceived as formal and task-oriented, can also foster a degree of social presence. Participants further described SAELM as proactive and motivated, with one noting, “Agent [SAELM] expressed enthusiasm and willingness to help,” as 1.2% of AAE comments were tagged as “enthusiastic.”
@@ 6. Discussion
Through a mixed-methods study combining quantitative and qualitative approaches, we investigated two sociolect-specific agents, namely AAELM and QSLM, representing AAE and Queer slang, alongside a SAE baseline. We found that AAE speakers rely on and prefer an LLM that communicates in SAE rather than AAE, primarily due to the LLM’s inability to use AAE in a natural or respectful manner. In contrast, Queer slang participants did not exhibit a strong preference for either the SAELM or QSLM, with user reliance on the LLM primarily correlating with increased social presence and reduced frustration when interacting with the QSLM.
Our study shows how increased reliance on an LLM is assocciated with increased user trust and increased perceived legitimacy of AI-generated outputs. We also observed how generally, even though users preferred and relied more on the LLM without anthropomorphic cues, this LLM also is less likely to be perceived as warm and authentic to the user. Finally, we observed how the degree to which users prefer an LLM with or without anthropomorphic cues also depends on the specific anthropomorphic cues present, as we see a difference in explicit preference for an LLM and certain perception features that differ between the studies that use AAE and Queer slang. Below we discuss the implications of our results and key takeaways from our variable findings.
@@
Our major takeaway was that AAE speakers overall relied on and preferred the SAE agent over the AAE agent. This is noteworthy as it shows personalization and anthropomorphic design of agents in this scenario can hinder not only user perception of an LLM, but also directly negatively affect the way they use it. These findings are further reinforced by our qualitative feedback, with AAE users describing the AAE agent’s responses as unnatural, abnormal, and disrespectful. Additionally, existing research highlights that over-anthropomorphized agents are often perceived as uncanny or even mocking (Gadiraju et al., 2023; Dhingra et al., 2023; Cheng et al., 2023). A possible explanation for AAE speaker greater reliance and preference for LLMs without anthropomorphic cues is the perception that mimicking AAE encroaches upon Black cultural spaces (Grieser, 2022). This idea is deeply tied to identity, where the use of a sociolect is often linked to the ability to self-identify with the associated cultural or social group (Pitts and Gallois, 2019). Since an LLM is not a person, it cannot meet this prerequisite. The issue, therefore, is not that the LLM appears ”too human,” but that it is perceived as overstepping societal boundaries and appropriating a linguistic identity it cannot authentically claim. We suggest integrating user-centered feedback to allow for real-time refinements in sociolect representation, ensuring alignment with user expectations and minimizing the risk of perceived inauthenticity or offensiveness.
@@
While reliance is correlated with preference and perception, our study demonstrates that these measures are not directly correlated. Notably, variability exists between variables such as reliance, preference, and social proximity, particularly for AAE speakers. For instance, while AAE speakers may prefer and rely on an SAE agent, this does not necessarily correspond to a closer sense of social proximity. These findings highlight the importance of measuring behavioral responses rather than relying solely on self-reported preferences and perceptions (Zhou et al., 2024).
Current studies in LLM design have focused solely on perceptions (Finch et al., 2025; Wang et al., 2023), this approach captures only part of the picture. Additionally, the dynamics between perception and behavior can become more nuanced and complex during extended use of LLMs (Long et al., 2024). As studies often rely on self-reported perceptions of LLMs (Cohn et al., 2024; Wu et al., 2024; Deshpande et al., 2023; Banovic et al., 2023; Jiang et al., 2023; Chen et al., 2024; Narayanan Venkit et al., 2023), users’ behaviors—such as their preferences and reliances—may tell a different story. Our work highlights the importance of capturing users’ preferences, perceptions, and behaviors comprehensively. We recommend robust design methodologies that account for these variables to better understand the nuanced relationships between them and provide a more holistic view of user interaction with LLMs.
@@
Our next major finding was that not all agents using a sociolect are relied on to the same degree. When comparing cross-study reliance rates between two parallel studies—one focused on AAE sociolect and the other on Queer slang—it is evident that the strong preference and reliance AAE users displayed for the SAE agent over their sociolect agent is not mirrored by Queer slang users. Queer slang participant sentiments further reflect this distinction, as Queer slang users provided nearly four times as many comments expressing positivity and nearly twice as many comments indicating resonance with the LLM.
@@ 7. Limitations and Future Work
In our experimental setup, participants interacted with previously generated suggestions provided by the LLM, where the correct answers to the questions were intentionally withheld. This design choice was made to prevent participants from reasoning about the plausibility of the suggested answers, focusing instead on the perceived quality of the LLM’s engagement (Zhou et al., 2025a).For LLM outputs specifically in non-standard dialects, other methods to produce higher quality and more realistic LLM generations for dialects, , particularly for the AAE and Queer slang dialects in our study,(Gupta et al., 2024; Zhou et al., 2025b), should be explored. While this setup effectively isolates the LLM’s ability to present sociolect-specific responses, it inherently limits the interaction to single-turn exchanges, which do not reflect the multi-turn conversational dynamics typically associated with LLM usage. The single-turn suggestion structure restricts participants from engaging in back-and-forth interaction with conversational agents, as well as avoiding the influence of answer accuracy on participant judgments. Future research should address these limitations by exploring more dynamic, multi-turn interactions that better reflect real-world conversational use cases of sociolect-infused LLMs.
@@
In our experimental setup, we examined two distinct sociolects— AAE) and Queer slang—in comparison to SAE. Although the experiment was not designed to directly compare the two sociolects against each other, we observed notable similarities in how users of these sociolects were perceived. These findings, however, are not generalizable to other sociolects. Future work should broaden the scope beyond these two cases to examine additional sociolects, particularly those with varying power dynamics relative to dominant forms of English, following approaches used in studies investigating dialectal differences in LLMs (e.g., Chicano English, Appalachian English, RP British English) (Faisal et al., 2024; Mire et al., 2025). Our findings suggest that the effects of sociolect usage are highly dependent on the context in which it appears. Given that our study examined sociolect usage exclusively within a question-and-answer format, future research should investigate how additional social and situational factors shape perceptions of acceptable sociolect use (Matthes, 2019).
@@ Appendix H Warmth Phrases and Confidence Expressions
Table 3. Table of 14 warmth phrases used in SAE suggestion generation.
@@ Appendix I AAELM In-Context Learning Prompt
After translating the 280 warmth-epistemic combinations into AAE, we conducted a verification process to assess which phrases were perceived as authentic AAE expressions. Leveraging a separate participant pool on Prolific, we administered sociolect verification surveys to identify the phrases most closely aligned with AAE (see Appendix §J for an excerpt of the verification survey). This participant pool was intentionally broader than the experimental group, as the goal was to gather diverse perspectives on which sociolect the phrases most closely aligned with. By opening verification to participants from all demographics, we aimed to ensure a comprehensive evaluation of the translated phrases. We presented groups of 20 phrases to batches of 5 participants at a time. Participants were tasked with identifying which sociolect they felt best aligned with the translated phrases. Only phrases that were deemed to represent AAE by at least 4 out of the 5 verifiers were selected for inclusion as AAE suggestions. This process was iterated until we achieved a diverse selection of 20 sociolect aligned phrases with unique confidence expressions. These finalized AAE translations were then assigned to the AAELM, while their original SAE versions were used for the SAELM in the AAELM experimental setup. At the outset, we randomized the assignment of sociolect translations to specific questions for both videos. To ensure consistency, the corresponding SAE translation was paired with the same question as its sociolect counterpart (see Appendix §L for question alignment).
@@ 3. How many shoes had white laces and how many shoes had black laces?
(a) Preferences for AAE and SAE.
(b) Preferences for Queer slang and SAE.
[…]
Two bar charts showing participant preferences for language models using either Standard American English or a sociolect. The left plot shows more participants selected the SAE agent over the African American English agent. The right plot shows near-equal preferences between the Queer slang and SAE agents.
AAE Queer Slang
mean_sample 6.992481e-01 0.504926
p-value 1.123979e-16 0.842924
d-value 4.339398e-01 0.00984
Table 15. RQ2: T-test using Cohen’s D 1-sample (with 0.5)
[…]
@@
Table 16. RQ2: T-tests with respective perception variables.

### v2

@@ 5.4.3. Negative Perceptions of the Sociolect LLMs
These positive sentiments were contrasted by a notable subset of participants who perceived the sociolect LLM’s use of sociolect as forced or artificial. Among Queer participants, 8.2% of comments were tagged with “negative emotions,” 6.98% with “exaggerated usage,” and 5.33% with “disrespectful.” Some participants expressed frustration at the QSLM’s overuse of slang, with one noting, “Even people who use LGBTQ slang don’t talk like that constantly. It would be annoying to have an AI constantly use slang phrases.” Similarly, AAE participants criticized the AAELM, describing its language as unnatural and inauthentic. Comments such as “Agent [AAELM] using AAE sounds like a joke and not natural” were reflected in 3.41% of comments tagged as “unnatural,” 2.61% tagged as “unserious,” and 1.4% tagged as “disrespectful.”
Participants also raised concerns about stereotyping and cultural insensitivity. AAE participants expressed discomfort, perceiving the AAELM’s use of AAE as stereotyping or mocking Black culture. One respondent noted, “I don’t like the idea of AI using such language, especially if it is being programmed by someone who isn’t Black.” Similarly, Queer slang participants criticized the QSLM, feeling that its use of Queer slang bordered on mockery, with comments such as, “The [QSLM] seems to be making a mockery of lgbt [sic].” These reactions highlight the risks of sociolect adaptation, especially when it involves sensitive cultural or linguistic contexts, potentially alienating users if the sociolect is perceived as inauthentic or disrespectful.
@@ 5.4.4. Positive Perceptions of the SAE-using LLMs
AAE participants often emphasized that SAELM met their expectations for how an AI should behave. Statements like “Agent [SAELM] sounds normal and like an actual AI” underscore this sentiment, with 3.8% of AAE comments tagged as “normal.” These responses suggest that SAE’s alignment with participants’ expectations of AI fosters a sense of credibility and trustworthiness.
While both participant groups frequently perceived SAELM as more relatable and human-like, some responses also indicated that it could connect on a personal level. For example, 1.4% of AAE participant comments were tagged as “resonates.” Remarks such as “I have a personal connection with him” reveal that SAE, while perceived as formal and task-oriented, can also foster a degree of social presence. Participants further described SAELM as proactive and motivated, with one noting, “Agent [SAELM] expressed enthusiasm and willingness to help,” as 1.2% of AAE comments were tagged as “enthusiastic.”
@@ 6. Discussion
Through a mixed-methods study combining quantitative and qualitative approaches, we investigated two sociolect-specific agents, namely AAELM and QSLM, representing AAE and Queer slang, alongside a SAE baseline. We found that AAE speakers rely on and prefer a LLM that communicates in SAE rather than AAE, primarily due to the LLM’s inability to use AAE in a natural or respectful manner. In contrast, Queer slang participants did not exhibit a strong preference for either the SAELM or QSLM, with user reliance on the LLM primarily correlating with increased social presence and reduced frustration when interacting with the QSLM.
Our study shows how increased reliance on a LLM is associated with increased user trust and increased perceived legitimacy of AI-generated outputs. We also observed how generally, even though users preferred and relied more on the LLM without anthropomorphic cues, this LLM also is less likely to be perceived as warm and authentic to the user. Finally, we observed how the degree to which users prefer a LLM with or without anthropomorphic cues also depends on the specific anthropomorphic cues present, as we see a difference in explicit preference for a LLM and certain perception features that differ between the studies that use AAE and Queer slang. Below we discuss the implications of our results and key takeaways from our variable findings.
@@
Our major takeaway was that AAE speakers overall relied on and preferred the SAELM over the AAELM. This is noteworthy as it shows personalization and anthropomorphic design of agents in this scenario can hinder not only user perception of a LLM, but also directly negatively affect the way they use it. These findings are further reinforced by our qualitative feedback, with AAE users describing the AAELM’s responses as unnatural, abnormal, and disrespectful. Additionally, existing research highlights that over-anthropomorphized agents are often perceived as uncanny or even mocking (Gadiraju et al., 2023; Dhingra et al., 2023; Cheng et al., 2023). A possible explanation for AAE speaker greater reliance and preference for LLMs without anthropomorphic cues is the perception that mimicking AAE encroaches upon Black cultural spaces (Grieser, 2022). This idea is deeply tied to identity, where the use of a sociolect is often linked to the ability to self-identify with the associated cultural or social group (Pitts and Gallois, 2019). Since a LLM is not a person, it cannot meet this prerequisite. The issue, therefore, is not that the LLM appears ”too human,” but that it is perceived as overstepping societal boundaries and appropriating a linguistic identity it cannot authentically claim. We suggest integrating user-centered feedback to allow for real-time refinements in sociolect representation, ensuring alignment with user expectations and minimizing the risk of perceived inauthenticity or offensiveness.
@@
While reliance is correlated with preference and perception, our study demonstrates that these measures are not directly correlated. Notably, variability exists between variables such as reliance, preference, and social proximity, particularly for AAE speakers. For instance, while AAE speakers may prefer and rely on an SAELM, this does not necessarily correspond to a closer sense of social proximity. These findings highlight the importance of measuring behavioral responses rather than relying solely on self-reported preferences and perceptions (Zhou et al., 2024).
Current studies in LLM design have focused solely on perceptions (Finch et al., 2025; Wang et al., 2023); this approach captures only part of the picture. Additionally, the dynamics between perception and behavior can become more nuanced and complex during extended use of LLMs (Long et al., 2024). As studies often rely on self-reported perceptions of LLMs (Cohn et al., 2024; Wu et al., 2024; Deshpande et al., 2023; Banovic et al., 2023; Jiang et al., 2023; Chen et al., 2024; Narayanan Venkit et al., 2023), users’ behaviors—such as their preferences and reliances—may tell a different story. Our work highlights the importance of capturing users’ preferences, perceptions, and behaviors comprehensively. We recommend robust design methodologies that account for these variables to better understand the nuanced relationships between them and provide a more holistic view of user interaction with LLMs.
@@
Our next major finding was that not all agents using a sociolect are relied on to the same degree. When comparing cross-study reliance rates between two parallel studies—one focused on AAE sociolect and the other on Queer slang—it is evident that the strong preference and reliance AAE users displayed for the SAELM over their sociolect agent is not mirrored by Queer slang users. Queer slang participant sentiments further reflect this distinction, as Queer slang users provided nearly four times as many comments expressing positivity and nearly twice as many comments indicating resonance with the LLM.
@@ 7. Limitations and Future Work
In our experimental setup, participants interacted with previously generated suggestions provided by the LLM, where the correct answers to the questions were intentionally withheld. This design choice was made to prevent participants from reasoning about the plausibility of the suggested answers, focusing instead on the perceived quality of the LLM’s engagement (Zhou et al., 2025a). For LLM outputs specifically in non-standard dialects, other methods to produce higher quality and more realistic LLM generations for dialects, particularly for the AAE and Queer slang dialects in our study, (Gupta et al., 2024; Zhou et al., 2025b), should be explored. While this setup effectively isolates the LLM’s ability to present sociolect-specific responses, it inherently limits the interaction to single-turn exchanges, which do not reflect the multi-turn conversational dynamics typically associated with LLM usage. The single-turn suggestion structure restricts participants from engaging in back-and-forth interaction with conversational agents, as well as avoiding the influence of answer accuracy on participant judgments. Future research should address these limitations by exploring more dynamic, multi-turn interactions that better reflect real-world conversational use cases of sociolect-infused LLMs.
@@
In our experimental setup, we examined two distinct sociolects— AAE and Queer slang— in comparison to SAE. Although the experiment was not designed to directly compare the two sociolects against each other, we observed notable similarities in how users of these sociolects were perceived. These findings, however, are not generalizable to other sociolects. Future work should broaden the scope beyond these two cases to examine additional sociolects, particularly those with varying power dynamics relative to dominant forms of English, following approaches used in studies investigating dialectal differences in LLMs (e.g., Chicano English, Appalachian English, RP British English) (Faisal et al., 2024; Mire et al., 2025).
Our findings suggest that the effects of sociolect usage are highly dependent on the context in which it appears. Given that our study examined sociolect usage exclusively within a question-and-answer format, future research should investigate how additional social and situational factors shape perceptions of acceptable sociolect use (Matthes, 2019).
@@ Appendix H Warmth Phrases and Confidence Expressions
Table 3. Warmth phrases used in SAE suggestion generation.
@@ Appendix I AAELM In-Context Learning Prompt
After translating the 196 warmth-epistemic combinations into AAE, we conducted a verification process to assess which phrases were perceived as authentic AAE expressions. Leveraging a separate participant pool on Prolific, we administered sociolect verification surveys to identify the phrases most closely aligned with AAE (see Appendix §J for an excerpt of the verification survey). This participant pool was intentionally broader than the experimental group, as the goal was to gather diverse perspectives on which sociolect the phrases most closely aligned with. By opening verification to participants from all demographics, we aimed to ensure a comprehensive evaluation of the translated phrases. We presented groups of 20 phrases to batches of 5 participants at a time. Participants were tasked with identifying which sociolect they felt best aligned with the translated phrases. Only phrases that were deemed to represent AAE by at least 4 out of the 5 verifiers were selected for inclusion as AAE suggestions. This process was iterated until we achieved a diverse selection of 20 sociolect aligned phrases with unique confidence expressions. These finalized AAE translations were then assigned to the AAELM, while their original SAE versions were used for the SAELM in the AAELM experimental setup. At the outset, we randomized the assignment of sociolect translations to specific questions for both videos. To ensure consistency, the corresponding SAE translation was paired with the same question as its sociolect counterpart (see Appendix §L for question alignment).
@@ 3. How many shoes had white laces and how many shoes had black laces?
Table 15. RQ2: T-tests with respective perception variables
@@
(a) Preferences for AAE and SAE
(b) Preferences for Queer slang and SAE
[…]
Two bar charts showing participant preferences for language models using either Standard American English or a sociolect. The left plot shows more participants selected the SAELM over the African American English agent. The right plot shows near-equal preferences between the Queer slang and SAELMs.
AAE Queer Slang
mean_sample 6.992481e-01 0.504926
p-value 1.123979e-16 0.842924
d-value 4.339398e-01 0.00984
Table 16. RQ2: T-test using Cohen’s D 1-sample (with 0.5)
[…]